OpenStudy (anonymous):

Is 3/4 bigger, equal to, or smaller than 4/7?

OpenStudy (anonymous):

3/4 = 0.75 4/8 = 3/4 = 0.75 4/8 is smaller smaller than 4/7 Meaning that 4/7 is greater than 3/4 Do you follow that path of logic?

OpenStudy (anonymous):

Oh, yes! I think I've got it, thank you. (Fractions are aaaawful. Lol)

OpenStudy (anonymous):

3/4=0.75 4/7=1.25 therefore 4/7 is greater.

OpenStudy (anonymous):

oops not 1.25

OpenStudy (anonymous):

0.571428571